1.A man deposits P5000 in a savings bank. If an interest of 7% is added to the principal at the end of each year, how much is his account at the end of 10years?
2.The value of a car at the end of each years is 2/3 of its value at the beginning of the year. Find the value of P385000 car after 5 years.
3.The population of a certain city is 950000. If the rate of increase is 10% each year, what will the population be at the end of 5 years?
4.A ball is thrown vertically upward to a height of 50 feet. Every time the ball strikes the ground, it rebounds 1/2 of the previous height. How far has the ball traveled up and down when it strikes the ground 6th time?

Answer by psbhowmick(878)

1. Compound interest problem:
A = P(1 + R/N)^NT
Here, P = 5000, N = 1, T = 10, R = 0.07
A = 5000(1+0.07)^10 = 9835.76
Try the rest yourself.